OpenStudy (anonymous):

Last math problem, need help need to find all 6 Regions

OpenStudy (math&ing001):

To find the elements in region I, we need to find the cardinalities of regions II and IV first. Can you do that ?

OpenStudy (anonymous):

I believe it's 6

OpenStudy (math&ing001):

For which region ?

OpenStudy (anonymous):

1

OpenStudy (math&ing001):

I got a slightly different result. Here's my reasoning: n(A^B)=13 so n(II)=13-4=9 n(A^C)=7 so n(IV)=7-4=3 n(I) = n(A) - n(II) - n(IV) - n(V) = 23 - 16 = 7

OpenStudy (anonymous):

I posted the wrong question. here it is

OpenStudy (math&ing001):

For this one : n(A^B)=12 so n(II)=12-4=8 n(A^C)=10 so n(IV)=10-4=6 n(I) = n(A) - n(II) - n(IV) - n(V) = 23 - 18 = 5